OpenStudy (anonymous):

log_1 3(x+5) + log_1 3(x-5) = 2

OpenStudy (anonymous):

\[\log_13(x+5) + \log_13(x-5) = 2\] solve for x:

OpenStudy (anonymous):

I would really appreciate if anyone could teach me how to do this question, struggling very hard.

OpenStudy (anonymous):

apply logx+logy=logxy

OpenStudy (anonymous):

it will come log-1(9*(x^2-25)=2 => x^2-25=1/9 now solve it for x

OpenStudy (anonymous):

Yup...... so... \[\log_{1} 3(x-5)*3(x+5) = 2\]

OpenStudy (anonymous):

\[\log_{1}9(x^2-25) = 2 \] \[9(x^2-25) = 1^2\] \[x^2-25 = \frac{ 1 }{ 9 }\] \[x^2=25+\frac{1 }{ 9 }\]

OpenStudy (anonymous):

Nw just Solve for x